# Learning Dexterous Manipulation Skills from Imperfect Simulations

### v1

@@ I Introduction
Reinforcement learning (RL) paired with sim-to-real transfer has recently delivered a number of promising results in dexterous manipulation [openai2018learning, openai2019solving, handa2023dextreme, chen2023visual, qi2022hand]. Policies trained in massively parallel simulation [makoviychuk2021isaac] with domain randomization [peng2018sim] have demonstrated strong robustness and generalization capabilities in the real world.
However, in practice, sim-to-real transfer faces two major limitations. First, due to the complexity of physics simulation, only a limited range of tasks can be accurately modeled. Prior work either relies on specialized techniques for high-fidelity simulation [narang2022factory, heiden2021disect] or seeks generalization through domain randomization [lee2020learning, kumar2021rma, openai2018learning, loquercio2021learning]. However, as the task becomes more dynamic, the sim-to-real gap grows [wang2024lessons], and simulation alone becomes insufficient. Second, existing sensing modalities have an intrinsic sim-to-real gap. While vision can be partially mitigated through domain randomization [singh2024dextrah], tactile sensing remains difficult to approximate reliably. Although some work aims to improve tactile simulation [wang2022tacto, akinola2025tacsl] or trains policies using alternative proxy representations [qi2023general, yang2024anyrotate, yin2025learning], these approaches cannot leverage the full power of tactile sensing. These limitations remain widely viewed as major constraints on the complexity of tasks that can be achieved.
On the other hand, teleoperation and imitation learning [chi2025diffusion, zhao2023learning] remove the need for simulation entirely. In this setting, policies can learn directly from real-world interactions and sensorimotor signals, which avoids the challenges introduced by sim-to-real transfer. However, teleoperating dexterous hands is challenging because of the intrinsic morphology differences between human and robot hands [arunachalam2022dexterous, arunachalam2023holo, qin2023anyteleop]. As a result, it is difficult to collect datasets that are large and diverse enough to achieve the desired behavior and generalization.
@@ II Related Work
Dexterous manipulation has been a long-standing challenge in robotics [bicchi2002hands, okamura2000overview]. Early work focused on classical model-based control and analytic grasp planning [rus1999hand, mordatch2012contact, fearing1986implementing, morgan2022complex, patidar2023hand]. Recent years have seen rapid progress in learning-based approaches, which can be grouped into two primary directions: sim-to-real learning paired with reinforcement learning [openai2018learning, qi2022hand, chen2023visual] and imitation learning from teleoperation [lin2024learning, arunachalam2023holo, qin2023anyteleop] or human data [wang2024dexcap, xu2025dexumi].
Both directions, however, face notable limitations. Sim-to-real methods benefit from large-scale simulated data and can generalize across diverse objects, yet they remain limited by inaccuracies in modeling complex contact dynamics and sensing, a challenge that becomes more significant as task complexity increases [wang2024lessons, huang2023dynamic]. Imitation learning benefits from multisensory real-world data, yet collecting high-quality dexterous demonstrations is considerably more difficult than collecting data for simpler end-effectors. Our work seeks to combine the strengths of both approaches. We use large-scale simulation to learn motion primitives while leveraging real-world data to close the dynamics and sensing gaps. Moreover, our skill-based framework enables efficient collection of dexterous real-world data by using the simulation-trained policy itself as a reusable skill primitive.
In the context of nut fastening and screwdriving, there has been recent work combining sim-to-real transfer with teleoperation. For example, Liu et al. [liu2025dexndm] build a residual model from real-world interactions to compensate for the sim-to-real gap and achieve robust in-hand manipulation. Yin et al. [yin2025dexteritygen] use simulation-trained policies as stability controllers to enable complex manipulation skills. Both approaches can be integrated with teleoperated arm control to complete these tasks. However, they do not produce autonomous policies that incorporate tactile sensing. Another line of work is Kumar et al. [kumar2025diffusion], who demonstrate screwdriver turning by combining learning with trajectory optimization. Noseworthy et al. [noseworthy2025forge] present an autonomous sim-to-real policy but only show results with a parallel-jaw gripper and do not demonstrate regrasping.
Another way to address the sim-to-real gap is to refine the policies in the real world. For example, Transic [jiang2024transic] shows that sim-to-real policies can adapt to complex real-world dynamics with only a few human interventions as demonstrations, although the demonstrations are primarily performed with simple end-effectors. In contrast, we apply this idea to dexterous hands. Maddukuri et al. [maddukuri2025sim] show that co-training with both simulation and real-world data can reduce the gap and improve manipulation performance.
@@ III-A Training a Reinforcement Learning Policy in Simulation
Simplified Object Modeling. Our goal is to design a simulation environment that enables fast training and encourages the emergence of desired finger gaits for rotation. To achieve this, we construct a simplified simulated object (Figure 2) that captures the essence of rotational motion. The object consists of a fixed cylindrical base with a nut or handle attached via a revolute joint. This setup allows the policy to learn rotational motion efficiently without relying on expensive contact-rich simulations. A similar idea was explored in [lin2024twisting] to model bottle caps using a heuristic friction design. In contrast, we further simplify the model, since we can leverage real-world data to compensate for the resulting dynamics mismatch.
Specifically, for the nut-bolt task, we use a thick triangular shape as the training object (Figure 2A). The extra thickness is used to prevent the policy from learning suboptimal strategies that apply a large force from the bottom. The learned policy also discovers a high-clearance gait that transfers well to diverse real-world nuts such as hexagonal and cube-shaped nuts. For the screwdriver task, where the primary difficulty arises from slippage around the handle, we use spherical primitives to keep the learned behavior conservative (Figure 2B). This observation, that different training shapes lead to different rotational gaits, is also consistent with the findings in [qi2022hand]. Note that these objects do not need to be visually aligned with real world objects, as they are only used to learn the coarse motions used for real world data collection, as we discussed in Section III-B.
@@
Training Pipeline. Following [qi2023general, wang2024lessons], we first train an oracle policy and then distill it into a sensorimotor policy. The oracle policy $\bm{f}$ is trained with access to an embedding of privileged information [chen2020learning] $\bm{z}_{t}$. The sensorimotor policy operates without privileged sensing and instead conditions on a predicted embedding $\hat{\bm{z}}_{t}=\bm{\phi}(\bm{h}_{t})$ inferred from proprioceptive history $\bm{h}_{t}$ by a prediction module $\bm{\phi}$.
@@
Sensorimotor Policy Training. The sensorimotor policy receives proprioceptive states and a latent code $\hat{\bm{z}}_{t}=\bm{\phi}(\bm{h}_{t})$ inferred from a 30-timestep history. We train the policy using DAgger [ross2011reduction]: at each step, the sensorimotor policy acts in the environment, while the oracle policy provides target actions and ground-truth privileged embeddings. The training objective is
@@
where $\hat{a}^{\text{Hand}}_{t}$ denotes the actions produced by the sensorimotor policy. The embedding predictor $\phi$ is optimized using Adam [kingma2014adam] until convergence.
Randomization. We apply domain randomization during training to improve the robustness of the RL policy [peng2018sim]. Following [wang2024lessons], we randomize the nut/handle mass, center of mass, friction coefficient, size, and PD gains, and we also add observation noise. Detailed parameters are provided in the appendix.
@@ III-B Real World Data Collection with Learned Policy
To bridge this gap, we introduce a skill-based assisted teleoperation for real-world data collection. The core idea is to reuse the simulation-trained policy as a skill primitive for finger motion control. Instead of commanding every joint individually, the human operator controls only the wrist movement and decides when to activate the skill primitive (Figure 3). Wrist position is specified using the Quest VR [^1] https://www.meta.com/quest/products/quest-2 controller’s joystick. This approach is inspired by [lin2024learning], but we use a much finer-grained skill for data collection.
This framework offers several advantages. First, it delegates complex finger motions to a robust simulation-trained policy that generalizes across different objects, eliminating the need for humans to learn finger coordination under morphological differences. Second, using a joystick for arm control enables precise and intuitive wrist positioning. Finally, collecting data in the real world provides the multisensory observations necessary for these tasks.
@@ III-C Behavior Cloning with Multisensory Data
Neural Network Architecture. We use a feedforward network as the policy. The past $K$ timesteps of observations $(\bm{q}_{t-K+1:t},\bm{c}_{t-K+1:t})$ are concatenated into a single feature vector. Tactile signals are first flattened and passed through an MLP. The fused feature vector is then processed by an hourglass encoder [newell2016stacked], which outputs the action predictions. We also apply an action chunking strategy [zhao2023learning, chi2025diffusion], where the policy predicts a sequence of future actions $\hat{\bm{a}}_{t:t+H}$ rather than a single-step action. We use $K=5$ and $H=16$ unless otherwise noted.
@@
where $\hat{\bm{a}}_{t:t+H}$ is the action chunk predicted by $\pi_{\text{BC}}$, and $\bm{a}_{t:t+H}$ is the corresponding expert sequence. This objective encourages consistent predictions over the full horizon. We train the policy using Adam [kingma2014adam] for 200 epochs and normalize observations following [barreiros2025careful].
@@ IV-A Experiment Setup
Simulation. We train our policies in IsaacGym [makoviychuk2021isaac] using 8,192 parallel environments. Each environment contains a simulated XHand and the simplified object model described in Section III-A. The simulation runs at $200\text{\,}\mathrm{Hz}$, with control applied at $20\text{\,}\mathrm{Hz}$. Each episode lasts up to 800 simulation steps ($40\text{\,}\mathrm{s}$).
@@ IV-C Screwdriving Experiments
The baseline behavior cloning model already achieves a 69.20% progress ratio, which substantially outperforms expert replay. This phenomenon, where a behavior cloning policy surpasses the policy that generated the data, is consistent with filtered behavior cloning [kumar2022should]. In this approach, only successful trials are used for training. A similar effect has also been reported in [wang2024lessons].
@@ IV-E Simulation Ablations
In simulation, we verify the design choices used to train the screwdriving policies. We evaluate our two-stage training procedure against two alternatives: training without any privileged information, and using an asymmetric actor-critic [pinto2017asymmetric] architecture where the critic has access to full tactile information while the actor does not. The actor in this setting is directly deployable in the real world. The results are shown in Figure 6.
@@ V-B Reward
TABLE V: Domain Randomization Parameters. Object scale is discretely sampled from the specified set and multiplied by the base scale. Mass, center of mass, friction, restitution, and PD controller gains are uniformly sampled at environment initialization. Observation and action noise are sampled i.i.d. from Gaussian distributions at each timestep. Following [openai2018learning], we apply a random disturbance force with magnitude $2.0m$ (where $m$ is object mass) with probability 0.25 at each timestep.

### v2

@@ I Introduction
Reinforcement learning (RL) paired with sim-to-real transfer has recently delivered a number of promising results in dexterous manipulation [32, 31, 9, 6, 36]. Policies trained in massively parallel simulation [24] with domain randomization [34] have demonstrated strong robustness and generalization capabilities in the real world.
However, in practice, sim-to-real transfer faces two major limitations. First, due to the complexity of physics simulation, only a limited range of tasks can be accurately modeled. Prior work either relies on specialized techniques for high-fidelity simulation [27, 10] or seeks generalization through domain randomization [18, 16, 32, 22]. However, as the task becomes more dynamic, the sim-to-real gap grows [45], and simulation alone becomes insufficient. Second, existing sensing modalities have an intrinsic sim-to-real gap. While vision can be partially mitigated through domain randomization [42], tactile sensing remains difficult to approximate reliably. Although some work aims to improve tactile simulation [46, 1] or trains policies using alternative proxy representations [37, 48, 49], these approaches cannot leverage the full power of tactile sensing. These limitations remain widely viewed as major constraints on the complexity of tasks that can be achieved.
On the other hand, teleoperation and imitation learning [7, 52] remove the need for simulation entirely. In this setting, policies can learn directly from real-world interactions and sensorimotor signals, which avoids the challenges introduced by sim-to-real transfer. However, teleoperating dexterous hands is challenging because of the intrinsic morphology differences between human and robot hands [3, 2, 38]. As a result, it is difficult to collect datasets that are large and diverse enough to achieve the desired behavior and generalization.
@@ II Related Work
Dexterous manipulation has been a long-standing challenge in robotics [4, 30]. Early work focused on classical model-based control and analytic grasp planning [40, 25, 8, 26, 33]. Recent years have seen rapid progress in learning-based approaches, which can be grouped into two primary directions: sim-to-real learning paired with reinforcement learning [32, 36, 6] and imitation learning from teleoperation [20, 2, 38] or human data [44, 47].
Both directions, however, face notable limitations. Sim-to-real methods benefit from large-scale simulated data and can generalize across diverse objects, yet they remain limited by inaccuracies in modeling complex contact dynamics and sensing, a challenge that becomes more significant as task complexity increases [45, 12]. Imitation learning benefits from multisensory real-world data, yet collecting high-quality dexterous demonstrations is considerably more difficult than collecting data for simpler end-effectors. Our work seeks to combine the strengths of both approaches. We use large-scale simulation to learn motion primitives while leveraging real-world data to close the dynamics and sensing gaps. Moreover, our skill-based framework enables efficient collection of dexterous real-world data by using the simulation-trained policy itself as a reusable skill primitive.
In the context of nut fastening and screwdriving, there has been recent work combining sim-to-real transfer with teleoperation. For example, Liu et al. [21] build a residual model from real-world interactions to compensate for the sim-to-real gap and achieve robust in-hand manipulation. Yin et al. [50] use simulation-trained policies as stability controllers to enable complex manipulation skills. Both approaches can be integrated with teleoperated arm control to complete these tasks. However, they do not produce autonomous policies that incorporate tactile sensing. Another line of work is Kumar et al. [15], who demonstrate screwdriver turning by combining learning with trajectory optimization. Noseworthy et al. [29] present an autonomous sim-to-real policy but only show results with a parallel-jaw gripper and do not demonstrate regrasping.
Another way to address the sim-to-real gap is to refine the policies in the real world. For example, Transic [13] shows that sim-to-real policies can adapt to complex real-world dynamics with only a few human interventions as demonstrations, although the demonstrations are primarily performed with simple end-effectors. In contrast, we apply this idea to dexterous hands. Maddukuri et al. [23] show that co-training with both simulation and real-world data can reduce the gap and improve manipulation performance. RLPD [51] and Sparsh-X [11] share a similar philosophy of augmenting modalities through human-guided data collection; however, they are not motivated by the possibility of learning skills from imperfect simulations, as their policies are designed to be directly transferable.
@@ III-A Training a Reinforcement Learning Policy in Simulation
Simplified Object Modeling. Our goal is to design a simulation environment that enables fast training and encourages the emergence of desired finger gaits for rotation. To achieve this, we construct a simplified simulated object (Figure 2) that captures the essence of rotational motion. The object consists of a fixed cylindrical base with a nut or handle attached via a revolute joint. This setup allows the policy to learn rotational motion efficiently without relying on expensive contact-rich simulations. A similar idea was explored in [19] to model bottle caps using a heuristic friction design. In contrast, we further simplify the model, since we can leverage real-world data to compensate for the resulting dynamics mismatch.
Specifically, for the nut-bolt task, we use a thick triangular shape as the training object (Figure 2A). The extra thickness is used to prevent the policy from learning suboptimal strategies that apply a large force from the bottom. The learned policy also discovers a high-clearance gait that transfers well to diverse real-world nuts such as hexagonal and cube-shaped nuts. For the screwdriver task, where the primary difficulty arises from slippage around the handle, we use spherical primitives to keep the learned behavior conservative (Figure 2B). This observation, that different training shapes lead to different rotational gaits, is also consistent with the findings in [36]. Note that these objects do not need to be visually aligned with real world objects, as they are only used to learn the coarse motions used for real world data collection, as we discussed in Section III-B.
@@
Training Pipeline. Following [37, 45], we first train an oracle policy and then distill it into a sensorimotor policy. The oracle policy $\bm{f}$ is trained with access to an embedding of privileged information [5] $\bm{z}_{t}$. The sensorimotor policy operates without privileged sensing and instead conditions on a predicted embedding $\hat{\bm{z}}_{t}=\bm{\phi}(\bm{h}_{t})$ inferred from proprioceptive history $\bm{h}_{t}$ by a prediction module $\bm{\phi}$.
@@
Sensorimotor Policy Training. The sensorimotor policy receives proprioceptive states and a latent code $\hat{\bm{z}}_{t}=\bm{\phi}(\bm{h}_{t})$ inferred from a 30-timestep history. We train the policy using DAgger [39]: at each step, the sensorimotor policy acts in the environment, while the oracle policy provides target actions and ground-truth privileged embeddings. The training objective is
@@
where $\hat{a}^{\text{Hand}}_{t}$ denotes the actions produced by the sensorimotor policy. The embedding predictor $\phi$ is optimized using Adam [14] until convergence.
Randomization. We apply domain randomization during training to improve the robustness of the RL policy [34]. Following [45], we randomize the nut/handle mass, center of mass, friction coefficient, size, and PD gains, and we also add observation noise. Detailed parameters are provided in the appendix.
@@ III-B Real World Data Collection with Learned Policy
To bridge this gap, we introduce a skill-based assisted teleoperation for real-world data collection. The core idea is to reuse the simulation-trained policy as a skill primitive for finger motion control. Instead of commanding every joint individually, the human operator controls only the wrist movement and decides when to activate the skill primitive (Figure 3). Wrist position is specified using the Quest VR [^1] https://www.meta.com/quest/products/quest-2 controller’s joystick. This approach is inspired by [20], but we use a much finer-grained skill for data collection.
This framework offers several advantages. First, it delegates complex finger motions to a robust simulation-trained policy that generalizes across different objects, eliminating the need for humans to learn finger coordination under morphological differences. Second, using a joystick for arm control enables precise and intuitive wrist positioning. While arm motion is relatively straightforward to simulate, we intentionally exclude it from the initial simulation training stage, as learning the specific downward progression required for nut-bolt fastening would necessitate simulating thread-level mechanics. Finally, collecting data in the physical environment provides the multisensory observations, particularly tactile signals, that are essential for reliable task completion but remain difficult to approximate reliably in simulation.
@@ III-C Behavior Cloning with Multisensory Data
Neural Network Architecture. We use a feedforward network as the policy. The past $K$ timesteps of observations $(\bm{q}_{t-K+1:t},\bm{c}_{t-K+1:t})$ are concatenated into a single feature vector. Tactile signals are first flattened and passed through an MLP. The fused feature vector is then processed by an hourglass encoder [28], which outputs the action predictions. We also apply an action chunking strategy [52, 7], where the policy predicts a sequence of future actions $\hat{\bm{a}}_{t:t+H}$ rather than a single-step action. We use $K=5$ and $H=16$ unless otherwise noted.
@@
where $\hat{\bm{a}}_{t:t+H}$ is the action chunk predicted by $\pi_{\text{BC}}$, and $\bm{a}_{t:t+H}$ is the corresponding expert sequence. This objective encourages consistent predictions over the full horizon. We train the policy using Adam [14] for 200 epochs and normalize observations following [43].
@@ IV-A Experiment Setup
Simulation. We train our policies in IsaacGym [24] using 8,192 parallel environments. Each environment contains a simulated XHand and the simplified object model described in Section III-A. The simulation runs at $200\text{\,}\mathrm{Hz}$, with control applied at $20\text{\,}\mathrm{Hz}$. Each episode lasts up to 800 simulation steps ($40\text{\,}\mathrm{s}$).
@@ IV-C Screwdriving Experiments
The baseline behavior cloning model already achieves a 69.20% progress ratio, which substantially outperforms expert replay. This phenomenon, where a behavior cloning policy surpasses the policy that generated the data, is consistent with filtered behavior cloning [17]. In this approach, only successful trials are used for training. A similar effect has also been reported in [45].
@@ IV-E Simulation Ablations
In simulation, we verify the design choices used to train the screwdriving policies. We evaluate our two-stage training procedure against two alternatives: training without any privileged information, and using an asymmetric actor-critic [35] architecture where the critic has access to full tactile information while the actor does not. The actor in this setting is directly deployable in the real world. The results are shown in Figure 6.
@@ V-B Reward
TABLE V: Domain Randomization Parameters. Object scale is discretely sampled from the specified set and multiplied by the base scale. Mass, center of mass, friction, restitution, and PD controller gains are uniformly sampled at environment initialization. Observation and action noise are sampled i.i.d. from Gaussian distributions at each timestep. Following [32], we apply a random disturbance force with magnitude $2.0m$ (where $m$ is object mass) with probability 0.25 at each timestep.